# Mixed-wet percolation on a dual square lattice

### v1

@@ 5 Summary and Conclusion
We presented a percolation model in the context of two-phase flow in a mixed-wet porous media, where the porous medium is made of a random mixture of two types of grains with opposite wetting properties. The grains are modeled by ordinary site percolation, whereas the bonds in the dual lattice that appear between two adjacent occupied-unoccupied sites on the primal lattice represent the network of pores with no capillary forces where two fluids can flow without any capillary barrier. It has been found that spanning clusters of such zero capillary barrier bonds occur between the two critical thresholds, $p_{c}=0.4072$, the next-nearest neighbour site percolation threshold and $0.59271-p_{c}=$ 0.5927, the ordinary site percolation threshold. At these two thresholds, one needs to occupy the bonds on the dual lattice with a unit probability to have mixed wet percolation. The bond clusters in the dual lattice are found to be the perimeters of the site clusters on the primal lattice. A perimeter bond cluster consists of a number of closed loops connected by knots and is very different from an ordinary bond percolation cluster. Note that, at these points, the randomness of the bond clusters in the dual lattice comes from the randomness of the site clusters on the primal lattice. However, for $p_{c}<p<(1-p_{c})$, the bonds are placed with a certain probability $\omega$ and a critical probability $\omega_{c}<1$ for each $p$ is found at which the network percolates for the first time. It can be noted that the value of $\omega_{c}$ is the least at $p=1/2$. In this region, the bond clusters no longer consist of closed loops. Rather, they have dangling bonds. The bond density $\rho$ on the dual lattice is found to be a quadratic function of the site concentration $p$ on the primal lattice, $\rho=2p(1-p)$. Because of the non-linear nature of $\rho$, a scaling theory for the geometrical properties is developed in terms of $\Delta p$ instead of $\Delta\rho$ preserving the critical exponents. The wrapping probabilities provide the critical thresholds, and their derivatives provide the correlation length exponent $\nu$. The other critical exponents associated with the geometrical quantities are estimated by studying them varying $\Delta p$ in the limit $\Delta p\to 0$ as well as by finite-size scaling analysis. The values of the critical exponents indicate that the mixed wet percolation in the whole range $p_{c}<p<(1-p_{c})$ belongs to the same universality class as ordinary percolation. Thus, there exists a continuous line of percolation thresholds for mixed wet percolation between the two extreme critical points defined by $p_{c}$ and $(1-p_{c})$. It seems that the new geometries of the clusters in mixed wet percolation are not enough to change the percolation universality class.

### v2

@@ 5 Summary and Conclusion
We presented a percolation model in the context of two-phase flow in a mixed-wet porous medium, where the porous medium is made of a random mixture of two types of grains with opposite wetting properties. The grains are modeled by ordinary site percolation, whereas the bonds in the dual lattice that appear between two adjacent occupied-unoccupied sites on the primal lattice represent the network of pores with no capillary forces. It has been found that spanning clusters of such zero capillary barrier bonds occur between the two critical thresholds, $p_{c}=0.4072$, the next-nearest neighbour site percolation threshold and $1-p_{c}=0.5927$, the ordinary site percolation threshold. At these two thresholds, one needs to occupy the bonds on the dual lattice with unit probability to have mixed wet percolation. The bond clusters in the dual lattice are found to be the perimeters of the clusters of either the occupied sites or the unoccupied sites on the primal lattice. Though the perimeter bond clusters consist of a number of cluster hulls, exterior and interior, connected by knots and are very different from an ordinary bond percolation cluster, the critical property remains the same as that of OSP. Looking at the perimeter bond clusters, one would expect the critical behaviour to be that of a percolation hull rather than percolation clusters. Since at these points, the randomness and tortuosity of the bond clusters in the dual lattice come from the randomness and tortuosity of the site clusters on the primal lattice, and hence we achieved an unexpected result. Even after the introduction of randomness in the bond occupation in the region, $p_{c}<p<(1-p_{c})$, the critical behaviour remains the same. However, we found a continuous line of percolation thresholds for mixed wet percolation between the two extreme critical points defined by $p_{c}$ and $(1-p_{c})$.